OpenStudy (angel123):

PLZ HELP!! One summer, conservation officials caught and tagged otters in a river flood plain. Later in the summer, 50 otters were caught and 32 had already been tagged. The estimated size of the otter population was 200 for that location. a. How many otters did the conservation officials originally tag? b. If a group of 10 otters was trapped in the fall, what is the probability that 3 of them would be tagged?

OpenStudy (angel123):

@agent0smith can you please help me again

OpenStudy (agent0smith):

Maybe when i get back.

OpenStudy (angel123):

okay thank you

OpenStudy (agent0smith):

50 otters were caught and 32 had already been tagged so 32/50 So multiply that by 200. For b it's another binomial probability like your last Q.

OpenStudy (angel123):

for part b so just to make sure p=3/10=0.3 q=7/10=0.7 n = 10 r=3 ?? @agent0smith

OpenStudy (agent0smith):

No. See last post for the probability.